# Strong anomalous diffusion for free-ranging birds

### v1

@@ Abstract
Natural and anomalous diffusion are widely observed and used to explore causes and consequences of movement across organisms, resulting in extensive use of the mean and mean-squared displacement (MSD). Using high-resolution data from over 70 million localizations of young and adult free-ranging Barn Owls (Tyto alba), we demonstrate the necessity of a broad spectrum of displacement moments to characterize bird movement across scales. The mean and MSD — interchangeable with moments $q=1$ and 2 — are insufficient special cases. We reveal empirical strong anomalous diffusion as a nonlinear growth of displacement moments according to $|\bm{x}(t)|^{q}\right>\sim t^{\lambda(q)}$. The moment spectrum function $\lambda(q)$ displays piecewise linearity with a critical moment marking the crossover point between two scaling regimes, linked to a combination of age-specific behavioral modes. A critical timescale of five minutes marks an unexpected transition from a convex $\lambda_{s}(q)$ to a concave $\lambda_{\ell}(q)$, related to environmental and behavioral constraints. Using two stochastic models of varying ecological complexity, we demonstrate that strong anomalous diffusion may be widespread in animal movement, underscoring the importance of expanding analysis beyond the average and MSD.
@@ I Introduction
Currently, stochastic diffusion and anomalous diffusion models, see definitions below, describe the movement of particles, molecules, and animals in a wide range of physical, chemical, biological, and ecological systems [13, 14]. In the context of organism movement, diffusion models are used to explain how individuals disperse over space and time, affecting population dynamics, spatial distribution, and interaction with the environment [15, 16]. Diffusion is also relevant in ecological settings, where dispersal and other movement phenomena determine the structure and spread of the population, influence the dynamics of metapopulation, and underlie predator-prey interactions and the coexistence of competitors [17]. Understanding the diffusion processes that underlie the movement of organisms is also crucial for addressing applied ecological concerns related to habitat fragmentation, biodiversity conservation, species invasion, and the management of natural resources [18, 19, 20]. Furthermore, within movement ecology [21, 11], models such as the random walk are essential to link movement metrics such as mean-displacement and travel-distance to underlying biological and environmental drivers. By employing a range of models, from minimal parameter models [22, 23, 24] to more sophisticated, parameterized simulations that incorporate detailed ecological contexts and correlations [25, 26, 27], empirical field data can be parsed into meaningful patterns and predictions. These models often offer insights into the mechanisms that govern movement and predict how organisms respond to environmental changes [28, 29, 30], thus bridging between empirical observation and theoretical understanding of movement.
@@
Common observables in movement studies are mean displacement and MSD, both central to quantifying the diffusive movements of organisms [31, 32]. In ecological studies, MSD is used to characterize dispersal rates and movement strategies, offering insights into the underlying mechanisms of organism movement for both normal and anomalous diffusion processes [33, 10, 34, 35]. In normal diffusion, e.g., Pearson’s random walk [1], the Gaussian central limit theorem results in an MSD increasing linearly with time according to $\langle\bm{x}^{2}(t)\rangle\sim Dt$, where $\bm{x}(t)=\vec{x}(t)-\vec{x}(0)$ is the normalized position at time $t$, and $D$ is the diffusion coefficient. Although this is a useful model for many natural processes, in many practical scenarios, the statistical properties of particle motion deviate from the classical Gaussian behavior [9]. In these systems, with so-called anomalous diffusion, the MSD exhibits a behavior inconsistent with linear scaling with time according to $\langle\bm{x}^{2}(t)\rangle\sim t^{\alpha}$, with $\alpha$ being the diffusion exponent. Here, $\alpha=1$ corresponds to normal diffusion, $\alpha<1$ to subdiffusion with slow spread, and $\alpha>1$ to superdiffusion with faster-than-normal spread.
@@
In many studies, the MSD is used to characterize the motion, be it normal or anomalous. However, this may lead to a wrong conclusion, as the MSD is a meaningful characteristic for processed where the distribution scales like $(1/t^{\alpha/2})F(\bm{x}/t^{\alpha/2})$ (e.g., Gaussian processes). Such a behavior is called mono-scaling, as $\bm{x}$ scales with $t^{\alpha/2}$. In a more general setting, it has been shown by the work of Vulpiani and colleagues [36, 37, 38] that the MSD is only one point in the spectrum of exponents describing the growth of the moments of order $q$. Thus, a useful tool to analyze data is the moment spectrum function $\lambda(q)$ defined by
$|\vec{x}(t)-\vec{x}(0)|^{q}\right>\equiv\left<|\bm{x}|^{q}\right>\sim t^{\lambda(q)},$ (1)
@@
where $q$ is a positive real number, and for the special case of $q=2$ we recover the MSD. For scale-invariant processes, the scaling exponent is linear in $q$, e.g., for Brownian motion it follows $\lambda(q)=q/2$ and for fractional Brownian motion [39] $q=\alpha/2$ for all $q$. A nonlinear dependence of $\lambda(q)$ in $q$ is termed strong anomalous diffusion (SAND) [36] and entails that the MSD is not indicative of displacement at varying spatial scales.
@@
In this manuscript, we study and model the movement patterns of Barn Owls (Tyto alba) based on the displacement moments of their movement. In Sec. III we show empirical evidence for SAND with bi-linear scaling in the displacement moments, revealing that the MSD can be situated at the transition between scaling regimes, such that it does not represent either behavior accurately. Interestingly, SAND is found for both adult and young Barn Owls during the first few month after fledging (so-called fledglings) with a convex bi-linear $\lambda_{s}(q)$ at times shorter than 5 minutes. Contrary to the convex moment spectrum function at short times, we find a qualitatively different concave function $\lambda_{\ell}(q)$ at longer times. In Sec. IV we employ two stochastic models with varying degrees of ecological complexity to account for the bi-linear scaling and the transition from convex to concave moment spectrum functions. First, in Sec. IV.1 a bounded Lévy-walk (BLW) model is used as a zero-order model of the observed phenomena, coupling large displacements to bounded movement within a home range. Second, in Sec. IV.2 we use an ecologically-oriented multi-mode correlated velocity model [26, 52, 27, 29, 53] to show that the observed phenomena can be attributed to a mixture of two distinct ecological behaviors: local area-restricted searches and long-term commutes [54, 55]. We show that by parameterizing the multi-mode correlated velocity model from data we account for the observed phenomena, and for the similarities and differences between adults and fledglings. Finally, in Sec. V we discuss our results and establish the importance of accounting for a wide and continuous range of moments in movement studies.
@@ II.1 Tracking
Sixty Barn Owls were tracked in the Hula Valley, Israel (33.10N, 35.61E) between May and December 2018, using ATLAS (Advanced Tracking and Localization of Animals in Real-Life Systems), a reverse GPS system that localizes extremely light-weight, low-cost tags [56, 57, 11]. Each ATLAS tag transmits a distinct radio signal which is detected by a network of base-stations distributed in the study area. Tag localization is computed using nanosecond-scale differences in signal time-of-arrival to each station, allowing for real-time tracking and alleviating the need to retrieve tags or have power-consuming remote-download capabilities. The individuals tracking frequency was 0.125 and 0.25 Hz, for the fledglings and adults, respectively. Localization errors are reported as a $2\times 2$ covariance matrix per localization. In this study we omit localizations with variance $>50^{2}\;m^{2}$, defined in terms of the trace over the covariance matrix. Furthermore, we filtered out nights in which many localizations are missing ($>70\%$). Notably, in accordance with the typical error reported by the system ($\sigma\simeq 5$ m) [56], we assume $10$ m to be the noise limit in our measurements. While the noise is practically much smaller, this is treated as an upper limit for any significant results.
Eighteen adults, breeding in nest boxes [58], were tracked both during and post-breeding, and we used the hatching date to define the breeding season by the 90 days following hatching of nestlings, as for this period the nestlings still depend on their parents [59]. In addition, 42 fledglings were tracked up to five months after fledging. In total, our analyses incorporated high-quality data for 7,391 nights and over 70 million localizations. We limited the analyses to movement data collected during the nights, the activity hours of the Barn Owl. In Fig. 1a we display 30-night movement tracks of an adult female Barn Owl in the Hula Valley. Notably, we have checked that the maximum displacement – $\max[|\bm{x}|]$ – performed in a single night is $2.6\pm 1.2$ km and is similar to the values $2.45\pm 0.93$ km and $2.77\pm 0.25$ reported in Refs. [60] and [61] for different populations of Barn Owls in Israel.
@@ II.2 Segmentation
To separate commutes from searches, daily tracks were segmented by detecting distinct switching points separating the two modes. We used the Penalized Contrast Method suggested by Barraquand and Benhamou [62], a non-parametric method in which the initial number of segments is unknown and estimated by minimizing a penalized contrast function. First passage time (FPT) was used as the focal metric [62]. Each point was assigned an FPT outside a radius of $R_{s}$ and data was segmented such that points with similar FPT that were close in time were clustered together [63]. Data are then split into commuting and searches according to a threshold on the mean FPT chosen in accordance with the animal’s velocity. In our segmentation we choose $R_{s}=100$ m and a threshold of $50$ s, see Ref. [55] for additional details.
@@ II.3 Data analysis
Fits for all moments and for the moment spectrum functions below were performed using SciPy library’s curve-fit (nonlinear least squares method) in python 3.12. In accordance with the observed bi-linear trends we have fitted a piecewise linear function to the observed curves, measuring the slopes at low and high moments and the transition point between the two slopes. As an error estimate we took one standard deviation of the fitted parameter. Notably, as several empirical tracks we have missing data points we only include in our analysis time-series in which $>$90% of the points are present. For any time-series that has $>$90% of the points present, we fill each missing data point with a NaN (Not a Number), i.e., an empty placeholder which is naturally not included in the moment averaging. To verify that this choice does not affect the statistics, we checked that our results do not change when varying the 90% threshold between 70% and 95%. Furthermore, we checked that in the simulations, randomly replacing 10% of the data points in a simulated ensemble with NaNs does not significantly affect the results.
@@ III Results
Figure 1: Results for the adult Barn Owls. (a) Thirty-one nights of ATLAS relocations of a Female Barn Owl in the Hula Valley at a mostly constant frequency of 0.25 Hz. Tracks are split into observational time-frames of $T=4$ hours. (b) Log-log plot of the average absolute displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ with $\mathcal{R}_{0}=2.5$ km representing the average maximum displacement in a single night. Data points for moment orders $q=1,1.5,2,2.5,3$ are shown, see legend. Moments show distinct scaling regimes, with a transition at $\sim 5$ minutes. (c) Moment spectrum function $\lambda(q)$ at short times ($<5$ minutes, red triangles) exhibiting convex behavior. The dashed blue and orange lines represent the slopes of the bi-linear fit for low and high moments, respectively, see values in legends. Inset: $\lambda(q)$ for commuting flights alone (omitting area-restricted searches) as identified by segmenting the full tracks using behavioural change point analysis. (d) Moment spectrum function $\lambda(q)$ for long time intervals ($>5$ minutes, red triangles), exhibiting concave behavior. The dashed blue and orange lines represent the slopes of the bi-linear fit.
@@
Results for the adult Barn Owls are shown in Fig. 1. In Fig. 1b we show the averaged displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ for $q$ values between 1 and 3. Here, $\mathcal{R}_{0}=2.5$ km corresponds to the average maximum displacement that Barn Owls cover in a night, and $\mathcal{R}_{0}^{q}$ is used to normalize the data in Fig. 1b, plotting the moments in a unitless format. For each moment $q$, we have performed linear fits for both short and long times, and for all moments we observe a crossover between two regimes at $\sim 5$ minutes. For example, the MSD ($q=2$) in Fig. 1b follows $\langle|\bm{x}|^{2}(t)\rangle\sim t^{1.25}$ for $t<5$ minutes, entailing superdiffusive dynamics with faster-than-linear spread, and $\langle\bm{x}^{2}(t)\rangle\sim t^{0.86}$ for $2405<t<$ 240 minutes, entailing subdiffusive dynamics with slower-than-linear spread. For the fledglings (Fig. S2a [65]) we have similar superdiffusive scaling $\langle\bm{x}^{2}(t)\rangle\sim t^{1.37}$ for $t<5$ minutes, but diffusive scaling $\langle\bm{x}^{2}(t)\rangle\sim t^{1.04}$ at $2405<t<$ 240 minutes. Importantly, the empirical 5-minute time scale separating the two regimes in Fig. 1b is directly related to the birds’ behavior and finite home range; see discussion below.
@@
We repeated the calculations of the moments for $q\in[0,5]$ at 0.5 intervals to measure $\lambda(q)$ for both adults (Fig. 1) and fledglings (Fig. S2 [65]). In the following we focus on $\lambda(q)=\lambda_{s}(q)$ at short times, $t<5$ minutes, plotted in Fig. 1c (returning below to $\lambda(q)=\lambda_{\ell}(q)$ at $2405<t<$ 240 minutes). The convex behavior of the moment spectrum function in Fig. 1c is one of the main results of this manuscript, entailing that free-ranging Barn Owls exhibit SAND. Furthermore, as we observe in Figs. 1c and S2b [65] a bi-linear trend, we fit a piecewise linear function to the moment spectrum function $\lambda_{s}(q)$, measuring the slopes at low and high moments and the transition point between the two slopes. For both adults and fledglings the displacement moments can be described by
$|\bm{x}|^{q}\right>\sim t^{\lambda_{s}(q)}\quad,\;\lambda_{s}(q)/q\sim\begin{cases}s_{1},&\text{if }q<q_{s},\\ s_{2},&\text{if }q\geq q_{s},\end{cases}$ (2)
@@
where $s_{1}<s_{2}$ are the exponents, such that $\lambda_{s}(q)$ is convex, and $q_{s}$ is the critical moment of the piecewise scaling. For adults (Fig. 1c) we find $s_{1}=0.63\pm 0.06$ and $s_{2}=0.98\pm 0.03$ for low and high moments, respectively, with $q_{s}=2.14\pm 0.18$. A qualitatively similar – but quantitatively different – convex moment spectrum function is observed for the fledglings, with $s_{1}=0.48\pm 0.05$, $s_{2}=0.79\pm 0.03$ and $q_{s}=0.67\pm 0.06$ (Fig. S2b [65]). Here, the fledglings have lower exponents for both low and high moments with a significantly lower value of $q_{s}$. The exponent values are summarized in Table 1.
@@
Table 1: Fit results for $\lambda_{s}(q)$ and $\lambda_{\ell}(q)$, for adults, fledglings, Bounded Lévy Walk (BLW) model with $\alpha=1.6$ and bounding radius $R=2$ km, and the multi-mode model parameterized for adults and parameterized for fledglings. The errors for the adults and fledglings are given in the text, and we have checked that all other errors are $<10\%$.
A tempting explanation for the differences between adults and fledglings is that the former display ballistic-like flights while the latter do not fly in a straight path for prolonged periods. However, this interpretation is incorrect. To show this we analyze an ensemble of directed flights (commutes) between distant locations. To obtain these commutes, we used a behavioural change point analysis segmentation method [26] termed the Penalized Contrast Method [62], see Methods. The segmentation yields two different ensembles: an ensemble of commutes and an ensemble of area-restricted searches. Following this behavior-based segmentation, we repeat the above analysis for only commutes, resulting in a single scaling of $0.92$ for the adults (Fig. 1c, inset) and of $0.94$ for the fleglings (Fig. S2b, inset [65]), both suggesting near ballistic mono-scaling. Thus, ballistic commutes are similar between adults and fledglings and cannot account for the difference between their different exponents, see further discussion below.
In contrast to the convex moment spectrum function for $t<5$ minutes, at long times, $t>5$ minutes, we find a concave moment spectrum function for both adults and fledglings (Figs. 1d and S2c [65]). This transition between a convex and concave $\lambda(q)$ is another key finding of this manuscript. Here, the moments also follow bi-linear behavior:
$|\bm{x}|^{q}\right>\sim t^{\lambda_{\ell}(q)}\quad,\;\lambda_{\ell}(q)/q\sim\begin{cases}\ell_{1},&\text{if }q<q_{\ell},\\ \ell_{2},&\text{if }q\geq q_{\ell},\end{cases}$ (3)
with $\ell_{1}>\ell_{2}$, resulting in a concave $\lambda_{\ell}(q)$. Fitting a piecewise linear function to $\lambda_{\ell}(q)$ for adults (Fig. 1d) we find at lower moments $\ell_{1}=0.5\pm 0.04$, transitioning at $q_{\ell}=1.57\pm 0.12$ to $\ell_{2}=0.07\pm 0.02$. Notably, the low value of $\ell_{2}$ suggests that the long-time moment function nearly saturates at high moments. Similar transition between a convex and a concave function are observed for the fledglings (Fig. S2c [65]), with elevated values of $\ell_{1}=0.74\pm 0.04$ and $\ell_{2}=0.2\pm 0.02$, and a lower value of $q_{\ell}=1.19\pm 0.07$. The exponent values for adults and fledglings are summarized in Table 1.

### v2

@@ Abstract
Diffusion and anomalous diffusion are widely observed and used to study movement across organisms, resulting in extensive use of the mean and mean-squared displacement (MSD). However, these measures – corresponding to specific displacement moments – do not capture the full complexity of movement behavior. Using high-resolution data from over 70 million localizations of young and adult free-ranging Barn Owls (Tyto alba), we reveal strong anomalous diffusion as nonlinear growth of displacement moments. The moment spectrum function $\lambda_{t}(q)$ – defined by $\left<|\bm{x}(t)|^{q}\right>\sim t^{\lambda_{t}(q)}$ – displays piecewise linearity in $q$, with a critical moment marking the crossover between scaling regimes. This highlights the need of a broad spectrum of displacement moments to characterize movement, which we link to age-specific ecological drivers. Furthermore, a characteristic timescale of five minutes marks an unexpected transition from a convex to a concave $\lambda_{t}(q)$. Using two stochastic models – a bounded Lévy walk and a multi-mode behavioral model – we account for the observed phenomena, showing good agreement with data, relating age-specific behavioral states to environmentally confined movement, and demonstrating how Lévy walk-like patterns can arise from underlying behavioral structure. Finally, we discuss the ecological significance of our results, arguing that strong anomalous diffusion may be widespread in animal movement.
@@ I Introduction
Currently, diffusion and anomalous diffusion models, see definitions below, describe the movement of particles, molecules, and animals in a wide range of physical, chemical, biological, and ecological systems [13, 14]. In the context of organism movement, diffusion models are used to explain how individuals disperse over space and time, affecting population dynamics, spatial distribution, and interaction with the environment [15, 16]. Diffusion is also relevant in ecological settings, where dispersal and other movement phenomena determine the structure and spread of the population, influence the dynamics of metapopulation, and underlie predator-prey interactions and the coexistence of competitors [17]. Understanding the diffusion processes that underlie the movement of organisms is also crucial for addressing applied ecological concerns related to habitat fragmentation, biodiversity conservation, species invasion, and the management of natural resources [18, 19, 20]. Furthermore, within movement ecology [21, 11], models such as the random walk are essential to link movement metrics such as mean-displacement and travel-distance to underlying biological and environmental drivers. By employing a range of models, from minimal parameter models [22, 23, 24] to more sophisticated, parameterized simulations that incorporate detailed ecological contexts and correlations [25, 26, 27], empirical field data can be parsed into meaningful patterns and predictions. These models often offer insights into the mechanisms that govern movement and predict how organisms respond to environmental changes [28, 29, 30], thus bridging between empirical observation and theoretical understanding of movement.
@@
Common observables in movement studies are the mean displacement and the MSD, both central to quantifying the diffusive movements of organisms [31, 32]. In ecological studies, MSD is used to characterize dispersal rates and movement strategies, offering insights into the underlying mechanisms of organism movement for both normal and anomalous diffusion processes [33, 10, 34, 35]. In normal diffusion, e.g., Pearson’s random walk [1], the Gaussian central limit theorem results in an MSD increasing linearly with time according to $\langle\bm{x}^{2}(t)\rangle\sim 2dDt$, where $\bm{x}(t)=\vec{x}(t)-\vec{x}(0)$ is the position at time $t$, relative to the initial position, $d$ is the spatial dimension, and $D$ is the diffusion coefficient. Although this is a useful model for many natural processes, in many practical scenarios, the statistical properties of particle motion deviate from the classical Gaussian behavior [9]. In these systems, with so-called anomalous diffusion, the MSD exhibits a behavior inconsistent with linear scaling with time according to $\langle\bm{x}^{2}(t)\rangle\sim t^{\alpha}$, with $\alpha$ being the diffusion exponent. Here, $\alpha=1$ corresponds to normal diffusion, $\alpha<1$ to subdiffusion with slow spread, and $\alpha>1$ to superdiffusion with faster-than-normal spread.
@@
In many studies, the MSD is used to characterize the motion, be it normal or anomalous. However, this may lead to a wrong conclusion, as the MSD is a meaningful characteristic for processes where the distribution scales like $(1/t^{\alpha/2})F(\bm{x}/t^{\alpha/2})$ (e.g., Gaussian processes). Such a behavior is called mono-scaling, as $\bm{x}$ scales with $t^{\alpha/2}$. However, it has been shown by the work of Vulpiani and colleagues [36, 37, 38] that in a more general setting the MSD is only one point in the spectrum of exponents describing the growth of the moments of order $q$. Thus, a useful tool to analyze data is the moment spectrum function $\lambda_{t}(q)$ defined by
$\left<|\vec{x}(t)-\vec{x}(0)|^{q}\right>\equiv\left<|\bm{x}|^{q}\right>\sim t^{\lambda_{t}(q)},$ (1)
@@
where $q$ is a positive real number, and for the special case of $q=2$ we recover the MSD. For scale-invariant processes, the scaling exponent is time-independent and linear in $q$, e.g., for Brownian motion it follows $\lambda_{t}(q)=q/2$ and for fractional Brownian motion [39] $\lambda_{t}(q)=\alpha/2$ for all $q$. A nonlinear dependence of $\lambda_{t}(q)$ in $q$ is termed strong anomalous diffusion (SAND) [36] and entails that the MSD is not indicative of displacement at varying spatial scales.
@@
In this manuscript, we study and model the movement patterns of Barn Owls (Tyto alba) based on the displacement moments of their movement. In Sec. II, we describe the tracking, numerical, and analytical methods. Importantly, the same dataset was previously analyzed in Ref. [52], and is revisited here from a new theoretical and numerical perspective, applying the SAND framework and identifying movement scales not previously reported. In Sec. III we show empirical evidence for SAND with bi-linear scaling in the displacement moments, revealing that the MSD can be situated at the transition between scaling regimes, such that it does not represent either behavior accurately. Interestingly, SAND is found for both adult and young Barn Owls during the first few month after fledging (so-called fledglings) with a convex bi-linear $\lambda_{s}(q)$ at times shorter than 5 minutes. Contrary to the convex moment spectrum function at short times, we find a qualitatively different concave function $\lambda_{\ell}(q)$ at longer times. In Sec. IV we employ two stochastic models with varying degrees of ecological complexity to account for the bi-linear scaling and the transition from convex to concave moment spectrum functions. First, in Sec. IV.1 a bounded Lévy-walk (BLW) model is used as a zero-order model of the observed phenomena, coupling large displacements to bounded movement within a home range. Second, in Sec. IV.2 we use an ecologically-oriented multi-mode correlated velocity model [26, 53, 27, 29, 54] to show that the observed phenomena can be attributed to a mixture of two distinct ecological behaviors: local area-restricted searches and large-scale commutes [55, 52]. We show that by parameterizing the multi-mode correlated velocity model from data we account for the observed phenomena, and for the similarities and differences between adults and fledglings. In Sec. V we interpret the empirical SAND and scaling transitions in relation to age-specific foraging strategies and behavioral timescales [56], and discuss the generality of our results with respect to animal movement in light of the broader literature on scale-dependent movement [8, 55]. Finally, in Sec. VI we summarize and establish the importance of accounting for a wide and continuous range of moments in movement studies.
@@ II.1 Tracking
Sixty Barn Owls were tracked in the Hula Valley, Israel (33.10N, 35.61E) between May and December 2018, using ATLAS (Advanced Tracking and Localization of Animals in Real-Life Systems), a reverse GPS system that localizes extremely light-weight, low-cost tags [57, 58, 11]. Each ATLAS tag transmits a distinct radio signal which is detected by a network of base-stations distributed in the study area. Tag localization is computed using nanosecond-scale differences in signal time-of-arrival to each station, allowing for real-time tracking and alleviating the need to retrieve tags or have power-consuming remote-download capabilities. The individuals tracking frequency was 0.125 and 0.25 Hz, for the fledglings and adults, respectively. Localization errors are reported as a $2\times 2$ covariance matrix per localization. In this study we omit localizations with variance $50^{2}\;m^{2}$, defined in terms of the trace over the covariance matrix. Furthermore, we filtered out nights in which many localizations are missing ($70\%$). Notably, in accordance with the typical error reported by the system ($\sigma\simeq 5$ m) [57], we assume $10$ m to be the noise limit in our measurements. While the noise is practically much smaller, this is treated as an upper limit for any significant results.
Eighteen adults, breeding in nest boxes [59], were tracked both during and post-breeding, and we used the hatching date to define the breeding season by the 90 days following hatching of nestlings, as for this period the nestlings still depend on their parents [60]. In addition, 42 fledglings were tracked up to five months after fledging. In total, our analyses incorporated high-quality data for 7,391 nights and over 70 million localizations. We limited the analyses to movement data collected during the nights, the activity hours of the Barn Owl.
In Figs.1a and 1e, we present thirty-night movement tracks of an adult female and a fledgling, respectively, in the Hula Valley. The movement patterns exhibit both localized search behavior and long-distance commutes [52]. Additionally, the spatial concentration of activity suggests a distinct home range, while the fledgling appears to engage in more long-distance commutes than the adult female. We elaborate on these points below. Finally, we verified that the maximum nightly displacement – $\max[|\bm{x}|]$ – is $2.6\pm 1.2$ km, consistent with previously reported values of $2.45\pm 0.93$ km and $2.77\pm 0.25$ km from other Barn Owl populations in Israel [61, 62].
@@ II.2 Segmentation
To separate long-distance commutes from localized search behavior, daily tracks were segmented by detecting distinct switching points separating the two modes. We used the Penalized Contrast Method suggested by Barraquand and Benhamou [63], a non-parametric method in which the initial number of segments is unknown and estimated by minimizing a penalized contrast function. First passage time (FPT) was used as the focal metric [63]. Each point was assigned an FPT outside a radius of $R_{s}$ and data was segmented such that points with similar FPT that were close in time were clustered together [64]. Data are then split into commuting and searches according to a threshold on the mean FPT chosen in accordance with the animal’s velocity. In our segmentation we choose $R_{s}=100$ m and a threshold of $50$ s, see Ref. [52] for additional details.
@@ II.3 Data analysis
Fits for all moments and for the moment spectrum functions below were performed using SciPy library’s curve-fit (nonlinear least squares method) in python 3.12. In accordance with the observed bi-linear trends we have fitted a piecewise linear function to the observed curves, measuring the slopes at low and high moments and the transition point between the two slopes. As an error estimate we took one standard deviation of the fitted parameter. In general, across all panels shown below, the symbol sizes are larger than or comparable to the errors. Notably, in several empirical tracks we have missing data points; consequently, we only include in our analysis, time-series in which >>90% of the points are present. For any time-series that has >>90% of the points present, we fill each missing data point with a NaN (Not a Number), i.e., an empty placeholder which is naturally not included in the moment averaging. To verify that this choice does not affect the statistics, we checked that our results do not change when varying the 90% threshold between 70% and 95%. Furthermore, we checked that in the simulations, randomly replacing 10% of the data points in a simulated ensemble with NaNs does not significantly affect the results.
@@ III Results
Figure 1: Results for the adult (a-d) and fledgling (e-h) Barn Owls. (a,e) Thirty nights of ATLAS relocations of a (a) Female and (e) fledgling Barn Owl in the Hula Valley at a mostly constant frequency of 0.25 Hz. Tracks are split into observational time-frames of $T=4$ hours. (b,f) Log-log plot of the average absolute displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ with $\mathcal{R}_{0}=2.5$ km representing the average maximum displacement in a single night. Data points for moment orders $q=1,1.5,2,2.5,3$ are shown, see legend. We omit data for other moments in these panels, for clarity. Moments show distinct scaling regimes, with a transition at $\sim 5$ minutes. (c,g) Moment spectrum function $\lambda_{s}(q)$ at short times (<5<5 minutes, red triangles) exhibiting convex behavior. The dashed blue and orange lines represent the slopes of the bi-linear fit for low and high moments, respectively, see values in legends. Insets: $\lambda(q)$ for commuting flights alone (omitting area-restricted searches) as identified by segmenting the full tracks using behavioural change point analysis. The $q$ values along the horizontal axis in the inset are identical to those in the main panel and span the same range. In addition, the axis labels of the main panels apply also to the insets. (d,h) Moment spectrum function $\lambda_{\ell}(q)$ for long time intervals (>5>5 minutes, red triangles), exhibiting concave behavior. The dashed blue and orange lines represent the slopes of the bi-linear fit.
@@
Results for the Barn Owls are shown in Fig. 1. In Figs. 1b and 1f we show the averaged displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ for $q$ values between 1 and 3. Here, $\mathcal{R}_{0}=2.5$ km corresponds to the average maximum displacement that Barn Owls cover in a night (see Sec. II.1), and $\mathcal{R}_{0}^{q}$ is used to normalize the data, plotting the moments in a unitless format. For each moment $q$, we have performed linear fits for both short and long times, and for all moments we observe a crossover between two regimes at $\sim 5$ minutes. For example, the MSD ($q=2$) in Fig. 1b follows $\langle|\bm{x}|^{2}(t)\rangle\sim t^{1.25}$ for $t<5$ minutes, entailing superdiffusive dynamics with faster-than-linear spread, and $\langle\bm{x}^{2}(t)\rangle\sim t^{0.86}$ for $5<t<240$ minutes, entailing subdiffusive dynamics with slower-than-linear spread. For the fledglings (Fig. 1f) we have similar superdiffusive scaling $\langle\bm{x}^{2}(t)\rangle\sim t^{1.37}$ for $t<5$ minutes, but diffusive scaling $\langle\bm{x}^{2}(t)\rangle\sim t^{1.04}$ at $5<t<240$ minutes. Importantly, the empirical 5-minute time scale separating the two regimes in Fig. 1b is directly related to the birds’ behavior and finite home range; see Sec. V below.
@@
We repeated the calculations of the moments for $q\in[0,5]$ at 0.5 intervals to measure $\lambda_{t}(q)$ for both adults and fledglings. In the following we focus on $\lambda_{t}(q)=\lambda_{s}(q)$ at short times, $t<5$ minutes, plotted in Fig. 1c (returning below to $\lambda_{t}(q)=\lambda_{\ell}(q)$ at long times, $5<t<240$ minutes). The convex behavior of the moment spectrum function in Figs. 1c and 1g is one of the main results of this manuscript, entailing that free-ranging Barn Owls exhibit SAND. As we observe in these figures a bi-linear trend, we fit a piecewise linear function to the moment spectrum function $\lambda_{s}(q)$, measuring the slopes at low and high moments and the transition point between the two slopes. For both adults and fledglings the displacement moments can be described by
$\left<|\bm{x}|^{q}\right>\sim t^{\lambda_{s}(q)}\quad,\;\lambda_{s}(q)/q\sim\begin{cases}s_{1},&\text{if }q<q_{s},\\ s_{2},&\text{if }q\geq q_{s},\end{cases}$ (2)
@@
where $s_{1}<s_{2}$ are the exponents, such that $\lambda_{s}(q)$ is convex, and $q_{s}$ is the critical moment of the piecewise scaling. For adults (Fig. 1c) we find $s_{1}=0.63\pm 0.06$ and $s_{2}=0.98\pm 0.03$ for low and high moments, respectively, with $q_{s}=2.14\pm 0.18$. A qualitatively similar – but quantitatively different – convex moment spectrum function is observed for the fledglings, with $s_{1}=0.48\pm 0.05$, $s_{2}=0.79\pm 0.03$ and $q_{s}=0.67\pm 0.06$ (Fig. 1g). Here, the fledglings have lower exponents for both low and high moments with a significantly lower value of $q_{s}$. The exponent values are summarized in Table 1.
@@
Table 1: Fit results for $\lambda_{s}(q)$ and $\lambda_{\ell}(q)$, for adults, fledglings, Bounded Lévy Walk (BLW) model with $\alpha=1.6$ and bounding radius $R=2$ km, and the multi-mode model parameterized for adults and parameterized for fledglings. The errors for the adults and fledglings are given in the text, and we have checked that all other errors are $10\%$.
A tempting explanation for the differences between adults and fledglings is that the former display ballistic-like flights while the latter do not fly in a straight path for prolonged periods. However, this interpretation is incorrect. To show this we analyze an ensemble of directed flights (commutes) between distant locations. To obtain these commutes, we used a behavioural change point analysis segmentation method [26] termed the Penalized Contrast Method [63], see Sec. II.2. The segmentation yields two different ensembles: an ensemble of commutes and an ensemble of area-restricted searches. Following this behavior-based segmentation, we repeat the above analysis for only commutes, resulting in a single scaling of $0.92$ for the adults (Fig. 1c, inset) and of $0.94$ for the fleglings (Fig. 1g, inset), both suggesting near ballistic mono-scaling. Thus, ballistic commutes are similar between adults and fledglings and cannot account for the difference between their different exponents, see further discussion below.
In contrast to the convex moment spectrum function for $t<5$ minutes, at long times, $t>5$ minutes, we find a concave moment spectrum function for both adults and fledglings (Figs. 1d and 1h). This transition between a convex $\lambda_{s}(q)$ and concave $\lambda_{\ell}(q)$ is another key finding of this manuscript. Here, the moments also follow bi-linear behavior:
$\left<|\bm{x}|^{q}\right>\sim t^{\lambda_{\ell}(q)}\quad,\;\lambda_{\ell}(q)/q\sim\begin{cases}\ell_{1},&\text{if }q<q_{\ell},\\ \ell_{2},&\text{if }q\geq q_{\ell},\end{cases}$ (3)
with $\ell_{1}>\ell_{2}$, resulting in a concave $\lambda_{\ell}(q)$. Fitting a piecewise linear function to $\lambda_{\ell}(q)$ for adults (Fig. 1d) we find at lower moments $\ell_{1}=0.5\pm 0.04$, transitioning at $q_{\ell}=1.57\pm 0.12$ to $\ell_{2}=0.07\pm 0.02$. Notably, the low value of $\ell_{2}$ suggests that the long-time moment function nearly saturates at high moments. Similar transition between a convex and a concave function are observed for the fledglings (Fig. 1h), with elevated values of $\ell_{1}=0.74\pm 0.04$ and $\ell_{2}=0.2\pm 0.02$, and a lower value of $q_{\ell}=1.19\pm 0.07$. The exponent values for adults and fledglings are summarized in Table 1.